Container A has a total volume of exactly 3/4 quart. Container B can hold exactly 4/7 of a quart less than 6 quarts. What fraction of container B must be filled in order to hold the same volume as Container A?
Answer by abhinav75(8)

capacity of container A = 3/4 quarts
capacity of container B = 6 - 4/7 = (6*7 - 4)/7 = 38/7 quarts
the required fraction is (3/4)/(38/7)
= (3*7)/(38*4)
= 21/152
the fraction is 21/152